Given the function f(x)=4x/3x-4 find the domain
Answer by stanbon(75887)

f(x)=4x/3x-4 find the domain
----
f(x) = (4x)/(3x-4)
----
The denominator cannot be zero.
Solve: 3x-4 = 0
x = 4/3
-----
Domain: All Real Numbers except x = 4/3
=======================
Cheers,
Stan H.
==============